# Reward-Aware Proto-Representations in Reinforcement Learning

### v1

@@ 6.2 Option Discovery
Our approach, reward-aware covering eigenoptions (RACE), is similar to CEO but learns the DR instead of the SR, using its top eigenvector to define the intrinsic reward. The full algorithm and hyperparameters are in Appendix C.3. We compare RACE with CEO and a uniform random walk in the environments from Figure 1, using 100-step episodes with no terminal states. To evaluate exploration, we measure the percentage of states visited; to assess risk awareness, we track the average reward per time step. Our goal is to test whether RACE enables reward-aware exploration. The results are depicted in Figure 4, where each point represents the average performance of one hyperparameter setting over 10 seeds.
@@ 6.3 Count-Based Exploration
Table 1: Model-free count-based exploration. Values in thousands (×$\times 10^{3}$). 95% conf. intervals shown in parentheses.
@@ 6.4 Transfer Learning
The DR has been applied in the transfer learning setting to compute the optimal policy when the rewards at terminal states change [2, 27]. However, such computations require access to the transition dynamics of the environment. With our extension of the DR to state-action-dependent rewards and default features, it is now possible to compute the optimal policy without requiring access to the transition dynamics. We compare our approach with successor features [SFs; 1], which allow efficient transfer when the rewards of any states change. Note, however, that in this work we only consider the setting when only the rewards at terminal states change, since this is the setting that DFs can be applied to. The comparison of DFs and SFs under this setting does not serve as evidence that SFs is not effective for the setting it is designed for.
Figure 5: Left: Four rooms with multiple goals. Right: Cumulative return across new terminal reward configurations. Curves are averaged over 50 runs. The shaded area shows 95% conf. interval.
We perform experiments in the environment shown in Figure 5, consisting of four goal states. We describe our procedure for learning and using the SFs and DFs for transfer in Appendix C.5, as well as the features $\bm{\phi}$ used. Figure 5 shows the performance of the DFs, the SFs computed with respect to different numbers of policies, and the optimal performance obtained by training an optimal policy under each new terminal state reward configuration using Q-learning [40]. As previously mentioned in Section 5.1, in this setting, DFs is extremely effective because it directly computes the optimal policy for new terminal state reward configurations, while SFs only computes a policy that is at least as good as the policies used to compute the SFs. Note, however, that DFs computes an optimal policy in the linearly solvable MDP setting that balances the reward function and the cost of deviating from the default policy.
@@ 8 Conclusion
In this work, we advance the theoretical foundation of the default representation (DR), a reward-aware proto-representation. Our contributions include: (1) deriving dynamic programming and TD methods to learn the DR, (2) analyzing its eigenspectrum, (3) extending it to function approximation via default features, and (4) generalizing proto-representation to the maximum entropy RL setting. Empirically, we show that the DR, by incorporating reward structure, yields qualitatively different behavior than the SR across tasks such as reward shaping, option discovery, exploration, and transfer.
This work lays a foundation for the DR as a potential middle ground between reward-agnostic and risk-sensitive approaches. While we focused on the tabular case for clarity and to be able to evaluate the DR across a range of use cases where reward-agnostic proto-representations have been applied, restricting the scope to this setting remains a limitation of this work. Future work should explore the use of the DR in more complex settings, starting with adapting neural network methods developed for approximating the SR [e.g., 15, 10, 6].
@@
We thank Brett Daley for useful discussions and feedback. This research was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chair Program, and Alberta Innovates. It was also enabled in part by computational resources provided by the Digital Research Alliance of Canada.
@@ B.1.1 Derivation of Equation 7
Equation 7 states that the $(s,s^{\prime})$-th entry of the DR can be written as:
${\mathbf{Z}}(s,s^{\prime})=\sum_{\tau_{s\to s^{\prime}}}{\mathbf{P}}^{\pi_{d}}(\tau_{s\to s^{\prime}})\exp\big{(}r(\tau_{s\to s^{\prime}})/\lambda\big{)}.$ (18)
@@
Denote a trajectory from $s_{i}$ to $s_{j}$ by $\tau_{s_{i}\to s_{j}}$, the above equation can be written in the following more compact form:
${\mathbf{Z}}(s_{i},s_{j})=\sum_{\tau_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi_{d}}(\tau_{s_{i}\to s_{j}})\exp\big{(}r(\tau_{s_{i}\to s_{j}})/\lambda\big{)},$ (26)
where $\tau_{s_{i}\to s_{j}}$ is a trajectory from state $s_{i}$ to state $s_{j}$, ${\mathbf{P}}^{\pi_{d}}(\tau_{s_{i}\to s_{j}})$ is the probability of following trajectory $\tau_{s_{i}\to s_{j}}$ under the policy $\pi_{d}$, and $r(\tau_{s_{i}\to s_{j}})$ is the sum of rewards obtained in trajectory $\tau_{s_{i}\to s_{j}}$. ∎
@@ B.1.2 Derivation of Equation 2
Equation 2 can be derived in a similar manner as Equation 7. Equation 2 states that the $(s_{i},s_{j})$-th entry of the SR can be written as:
${\mathbf{\Psi}}^{\pi}(s_{i},s_{j})=\sum_{\tau_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi}(\tau_{s_{i}\to s_{j}})\gamma^{\eta(\tau_{s_{i}\to s_{j}})}.$ (27)
@@
${\mathbf{\Psi}}^{\pi}(s_{i},s_{j})=\sum_{\tau_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi}(\tau_{s_{i}\to s_{j}})\gamma^{\eta(\tau_{s_{i}\to s_{j}})},$ (36)
where $\tau_{s_{i}\to s_{j}}$ is a trajectory from state $s_{i}$ to state $s_{j}$, ${\mathbf{P}}^{\pi}(\tau_{s_{i}\to s_{j}})$ is the probability of following trajectory $\tau_{s_{i}\to s_{j}}$ under the policy $\pi$, and $\eta(\tau_{s_{i}\to s_{j}})$ is the number of steps it takes to visit $s_{j}$ from $s_{i}$ in trajectory $\tau_{s_{i}\to s_{j}}$. Note that when $\tau=(s_{i})=(s_{j})$, $\eta(\tau)=0$. ∎
@@
Suppose both the SR and DR are computed with respect to the same policy, i.e., $\pi=\pi_{d}$. When the reward function is constant, i.e., $r(s)=r(s^{\prime})\ \forall s,s^{\prime}\in\mathcal{S}$, the $i$-th eigenvectors of the SR and DR are equivalent, and the $i$-th eigenvalues of the SR ($\lambda_{\text{SR},i}$) and DR ($\lambda_{\text{DR},i}$) are related as follows:
$\lambda_{\text{SR},i}=\bigg{[}\gamma\Big{(}\lambda_{\text{DR},i}^{-1}-\exp\big{(}-r(s)/\lambda\big{)}+\gamma^{-1}\Big{)}\bigg{]}^{-1},$ (37)
where $\gamma$ is the discount factor of the SR, $r(s)$ is the state reward, and $\lambda$ is the relative importance of the deviation cost of the DR.
@@
Let $\lambda_{i},{\mathbf{e}}_{i}$ be the $i$-th eigenvalue and eigenvector of the DR. Assuming that the reward function is constant, and letting $\tilde{r}=\exp\big{(}-r(s)/\lambda\big{)}$, the DR can be written as $[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}$. Then, we have:
$[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}{\mathbf{e}}_{i}$ $=\lambda_{i}{\mathbf{e}}_{i}$ (38)
$[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]{\mathbf{e}}_{i}$ $=\lambda_{i}^{-1}{\mathbf{e}}_{i}$ (39)
$-\gamma{\mathbf{P}}^{\pi}{\mathbf{e}}_{i}$ $=\gamma(\lambda_{i}^{-1}-\tilde{r}){\mathbf{e}}_{i}$ (40)
$({\mathbf{I}}-\gamma{\mathbf{P}}^{\pi}){\mathbf{e}}_{i}$ $=\gamma(\lambda_{i}^{-1}-\tilde{r}+\gamma^{-1}){\mathbf{e}}_{i}$ (41)
$({\mathbf{I}}-\gamma{\mathbf{P}}^{\pi})^{-1}{\mathbf{e}}_{i}$ $=\big{[}\gamma(\lambda_{i}^{-1}-\tilde{r}+\gamma^{-1})\big{]}^{-1}{\mathbf{e}}_{i}$ (42)
${\mathbf{\Psi}}^{\pi}{\mathbf{e}}_{i}$ $=\bigg{[}\gamma\Big{(}\lambda_{i}^{-1}-\exp\big{(}-r(s)/\lambda\big{)}+\gamma^{-1}\Big{)}\bigg{]}^{-1}{\mathbf{e}}_{i}$ (43)
Note that when $\lambda_{i}$ increases, $\big{[}\gamma(\lambda_{i}^{-1}-\tilde{r}+\gamma^{-1})\big{]}^{-1}$ increases. Therefore, the $i$-th largest eigenvector of the SR is equivalent to the $i$-th largest eigenvector of the DR. ∎
@@ B.3 Extension of the DR to State-Action-Dependent Reward
In this subsection, we derive the DR for the case when the reward function depends on both the state and action. For this setting, instead of formulating the deviation cost as $\operatorname{KL}\big{(}p^{\pi}(\cdot|S_{t})\|p^{\pi_{d}}(\cdot|S_{t})\big{)}$, we formulate the deviation cost from the default policy as $\operatorname{KL}\big{(}p^{\pi}(\cdot,\cdot|S_{t},A_{t})\|p^{\pi_{d}}(\cdot,\cdot|S_{t},A_{t})\big{)}$, where $p^{\pi}(\cdot,\cdot|S_{t},A_{t})$ denotes the distribution over the next state-action pairs given the current state-action pair $(S_{t},A_{t})$ and the policy $\pi$. Then, at every time step $t$, the agent receives a reward of $\tilde{r}(S_{t},A_{t})=r(S_{t},A_{t})-\lambda\operatorname{KL}\big{(}p^{\pi}(\cdot,\cdot|S_{t},A_{t})\|p^{\pi_{d}}(\cdot,\cdot|S_{t},A_{t})\big{)}$, where $\lambda>0$ determines the relative importance of the deviation cost.

### v2

@@ 6.2 Option Discovery
Our approach, reward-aware covering eigenoptions (RACE), is similar to CEO but learns the DR instead of the SR, using its top eigenvector to define the intrinsic reward. In RACE, at every iteration, the agent collects samples to learn the DR, the eigenvector of which is then used to compute an eigenoption. The eigenoption will be used in the next RACE iteration to collect samples. This forms a cycle of using eigenoptions to improve the learned representation, and using the learned representation to then refine eigenoptions. The full algorithm and hyperparameters are in Appendix C.3. We compare RACE with CEO and a uniform random walk in the environments from Figure 1, using 100-step episodes with no terminal states. To evaluate exploration, we measure the percentage of states visited; to assess risk awareness, we track the average reward per time step. Our goal is to test whether RACE enables reward-aware exploration. The results are depicted in Figure 4, where each point represents the average performance of one hyperparameter setting over 10 seeds.
@@ 6.3 Count-Based Exploration
Table 1: Model-free count-based exploration. Values in thousands ($\times 10^{3}$). 95% conf. intervals shown in parentheses.
@@ 6.4 Transfer Learning
The DR has been applied in the transfer learning setting to compute the optimal policy when the rewards at terminal states change [3, 34]. However, such computations require access to the transition dynamics of the environment. With our extension of the DR to state-action-dependent rewards and default features, it is now possible to compute the optimal policy without requiring access to the transition dynamics. We compare our approach with successor features [SFs; 2], which allow efficient transfer when the rewards of any states change. Note, however, that in this work we only consider the setting when only the rewards at terminal states change, since this is the setting that DFs can be applied to. The comparison of DFs and SFs under this setting does not serve as evidence that SFs is not effective for the setting it is designed for.
Figure 5: Left: Four rooms with multiple goals. Right: Cumulative return across new terminal state reward configurations. Curves are averaged over 50 runs. The shaded area shows 95% conf. interval.
We perform experiments in the environment shown in Figure 5, consisting of four goal states. We describe our procedure for learning and using the SFs and DFs for transfer in Appendix C.5, as well as the features $\bm{\phi}$ used. Figure 5 shows the performance of the DFs, the SFs computed with respect to different numbers of policies, and the optimal performance obtained by training an optimal policy under each new terminal state reward configuration using Q-learning [51]. As previously mentioned in Section 5, in this setting, DFs is extremely effective because it directly computes the optimal policy for new terminal state reward configurations, while SFs only computes a policy that is at least as good as the policies used to compute the SFs. Note, however, that DFs computes an optimal policy in the linearly solvable MDP setting that balances the reward function and the cost of deviating from the default policy.
@@ 8 Conclusion
In this work, we advance the theoretical foundation of the default representation (DR), a reward-aware proto-representation. Our contributions include: (1) deriving dynamic programming and TD methods to learn the DR, (2) analyzing its eigenspectrum, and (3) extending it to function approximation via default features. Empirically, we show that the DR, by incorporating reward structure, yields qualitatively different behavior than the SR across tasks such as reward shaping, option discovery, exploration, and transfer.
While the DR provides benefits in environments where being reward-aware is important, there are some limitations. First, naively computing the DR and its eigendecomposition can lead to numerical instabilities due to taking the exponential of negative rewards, as discussed in Appendix C.1. This could be problematic, especially in more complicated environments with long horizons. Second, as we show in Figures 6 and 8 in Appendix C, in environments in which all non-terminal state rewards are the same, the DR will perform as well as the SR. Given the numerical issues associated with the DR, the SR may be a better choice in these environments. Third, while the DR captures the environment dynamics more fully, this comes at the expense of flexibility. While the SR, which is reward-agnostic, can allow transfer learning when any state rewards change, the DR, which is reward-aware, can only perform transfer in settings where only the terminal state rewards change.
Finally, this work lays a foundation for the DR as a potential middle ground between reward-agnostic and risk-sensitive approaches. While we focused on the tabular case for clarity and to be able to evaluate the DR across a range of use cases where reward-agnostic proto-representations have been applied, restricting the scope to this setting remains a limitation of this work. Future work should explore the use of the DR in more complex settings, starting with adapting neural network methods developed for approximating the SR [e.g., 19, 14, 9].
@@
We thank Brett Daley for useful discussions and feedback. We thank anonymous reviewers whose feedback improved this work. This research was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chair Program, and Alberta Innovates. It was also enabled in part by computational resources provided by the Digital Research Alliance of Canada.
@@ B.1.1 Derivation of Equation 7
Equation 7 states that the $(s_{i},s_{j})$-th entry of the DR can be written as:
${\mathbf{Z}}(s_{i},s_{j})=\sum_{\tau\in\mathcal{T}_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi_{d}}(\tau)\exp\big(r(\tau)/\lambda\big).$ (17)
@@
The above equation can be written in the following more compact form:
${\mathbf{Z}}(s_{i},s_{j})=\sum_{\tau\in\mathcal{T}_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi_{d}}(\tau)\exp\big(r(\tau)/\lambda\big),$ (25)
where $\mathcal{T}_{s_{i}\to s_{j}}$ is the set of trajectories from state $s_{i}$ to state $s_{j}$, ${\mathbf{P}}^{\pi_{d}}(\tau)$ is the probability of following trajectory $\tau$ under the policy $\pi_{d}$, and $r(\tau)$ is the sum of rewards obtained in trajectory $\tau$. ∎
@@ B.1.2 Derivation of Equation 2
Equation 2 appeared in the work by Blier et al. [5], but the derivation was not shown. For completeness, we provide the derivation here, which is similar to that for Equation 7. Equation 2 states that the $(s_{i},s_{j})$-th entry of the SR can be written as:
${\mathbf{\Psi}}^{\pi}(s_{i},s_{j})=\sum_{\tau\in\mathcal{T}_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi}(\tau)\gamma^{\eta(\tau)}.$ (26)
@@
${\mathbf{\Psi}}^{\pi}(s_{i},s_{j})=\sum_{\tau\in\mathcal{T}_{s_{i}\to s_{j}}}{\mathbf{P}}^{\pi}(\tau)\gamma^{\eta(\tau)},$ (35)
where $\mathcal{T}_{s_{i}\to s_{j}}$ is the set of trajectories from state $s_{i}$ to state $s_{j}$, ${\mathbf{P}}^{\pi}(\tau)$ is the probability of following trajectory $\tau$ under the policy $\pi$, and $\eta(\tau)$ is the number of steps it takes to visit $s_{j}$ from $s_{i}$ in trajectory $\tau$. Note that when $\tau=(s_{i})=(s_{j})$, $\eta(\tau)=0$. ∎
@@
Suppose both the SR and DR are computed with respect to the same policy, i.e., $\pi=\pi_{d}$. When the reward function is constant and negative, i.e., $r(s)=r(s^{\prime})<0\ \forall s,s^{\prime}\in\mathcal{S}$, the SR and DR share the same set of eigenvectors. Furthermore, when the SR and DR are symmetric, the $i$-th eigenvectors of the SR and DR are equivalent, and the $i$-th eigenvalues of the SR ($\mu_{\text{SR},i}$) and DR ($\mu_{\text{DR},i}$) are related as follows:
$\mu_{\text{SR},i}=\bigg[\gamma\Big(\mu_{\text{DR},i}^{-1}-\exp\big(-r(s)/\lambda\big)+\gamma^{-1}\Big)\bigg]^{-1},$ (36)
where $\gamma\in(0,1)$ is the discount factor of the SR, $r(s)$ is the state reward, and $\lambda$ is the relative importance of the deviation cost of the DR.
@@
Let $\mu_{i},{\mathbf{e}}_{i}$ be the $i$-th eigenvalue and eigenvector of the DR. Assuming that the reward function is constant, and letting $\tilde{r}=\exp\big(-r(s)/\lambda\big)$, the DR can be written as $\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}$. Then, we have:
$[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]^{-1}{\mathbf{e}}_{i}$ $=\mu_{i}{\mathbf{e}}_{i}$ (37)
$[\tilde{r}{\mathbf{I}}-{\mathbf{P}}^{\pi}]{\mathbf{e}}_{i}$ $=\mu_{i}^{-1}{\mathbf{e}}_{i}$ (38)
$-\gamma{\mathbf{P}}^{\pi}{\mathbf{e}}_{i}$ $=\gamma(\mu_{i}^{-1}-\tilde{r}){\mathbf{e}}_{i}$ (39)
$({\mathbf{I}}-\gamma{\mathbf{P}}^{\pi}){\mathbf{e}}_{i}$ $=\gamma(\mu_{i}^{-1}-\tilde{r}+\gamma^{-1}){\mathbf{e}}_{i}$ (40)
$({\mathbf{I}}-\gamma{\mathbf{P}}^{\pi})^{-1}{\mathbf{e}}_{i}$ $=\big[\gamma(\mu_{i}^{-1}-\tilde{r}+\gamma^{-1})\big]^{-1}{\mathbf{e}}_{i}$ (41)
${\mathbf{\Psi}}^{\pi}{\mathbf{e}}_{i}$ $=\bigg[\gamma\Big(\mu_{i}^{-1}-\exp\big(-r(s)/\lambda\big)+\gamma^{-1}\Big)\bigg]^{-1}{\mathbf{e}}_{i}$ (42)
We have thus shown that the SR and DR share the same set of eigenvectors. When the SR and DR are symmetric, their eigenvalues are real, and we can further show that the orders of the eigenvectors of the SR and DR by their corresponding eigenvalues are identical.
From Eq. 42, we know that the eigenvalues of the SR and DR are related by the following function:
$\mu_{SR}=f(\mu_{DR})=[\gamma(\mu_{DR}^{-1}-\exp(-r(s)/\lambda)+\gamma^{-1})]^{-1}$ (43)
Taking the derivative, we have
$f^{\prime}(\mu_{DR})={\gamma^{-1}[1+\mu_{DR}(\gamma^{-1}-\exp(-r(s)/\lambda))]^{-2}}.$ (44)
When $\gamma=\exp(r(s)/\lambda)$, $f^{\prime}(\mu_{DR})$ is always positive, so the orders of the eigenvectors of the SR and DR are identical. When $\gamma\neq\exp(r(s)/\lambda)$, $f^{\prime}(\mu_{DR})>0$ except at the vertical asymptote at $\mu_{DR}=\frac{-1}{\gamma^{-1}-\exp(-r(s)/\lambda)}$. We now show that all eigenvalues of the DR lie on the same side of the vertical asymptote, and so the orders of eigenvectors are still identical. We do so by first deriving lower and upper bounds on the eigenvalues of the DR.
First, we show that all eigenvalues of the DR are positive. We start with the inverse of the DR, $\mathbf{Z}^{-1}=\operatorname{diag}(\exp(-\mathbf{r}/\lambda))-\mathbf{P}^{\pi_{d}}$. Consider a row $i$ of $\mathbf{Z}^{-1}$ corresponding to a non-terminal state. We know that $\mathbf{Z}^{-1}(i,i)=\exp(-r(s)/\lambda)-\mathbf{P}^{\pi_{d}}(i,i)>1-\mathbf{P}^{\pi_{d}}(i,i)=\sum_{j\neq i}|\mathbf{Z}^{-1}(i,j)|$. For a row $i$ corresponding to a terminal state, $\mathbf{Z}^{-1}(i,i)>\sum_{j\neq i}|\mathbf{Z}^{-1}(i,j)|$ easily holds. Then, by the Gershgorin circle theorem, we know that this matrix has all positive eigenvalues, and thus, all eigenvalues of the DR are positive.
Next, we derive an upper bound on the eigenvalues of the DR. Also applying the Gershgorin circle theorem on the inverse of the DR, we know that the set of eigenvalues of the DR lies within the union of the Gershgorin discs. For the $i$-th row of the inverse of the DR, if it corresponds to a non-terminal state, we have a disc centered at $\exp(-r(s)/\lambda)-\mathbf{P}^{\pi_{d}}(i,i)$ with radius $1-\mathbf{P}^{\pi_{d}}(i,i)$. For a terminal state, the disc is centered at $\exp(-r(s)/\lambda)$ with radius 0. It is then not hard to see that the minimum of the union of the discs is simply $\exp(-r(s)/\lambda)-1$. Since the lower bound on the eigenvalues of the inverse of the DR is $\exp(-r(s)/\lambda)-1$, the upper bound on the eigenvalues of the DR is $\frac{1}{\exp(-r(s)/\lambda)-1}$.
We now know that all eigenvalues of the DR lie in the range $\left(0,\frac{1}{\exp(-r(s)/\lambda)-1}\right]$. We now show that the vertical asymptote of $f$, if it exists, always lies outside of this range. We focus on the following two cases: 1) $\gamma<\exp(r(s)/\lambda)$, and 2) $\gamma>\exp(r(s)/\lambda)$.
First, when $\gamma<\exp(r(s)/\lambda)$, it can be easily shown that the vertical asymptote $\frac{-1}{\gamma^{-1}-\exp(-r(s)/\lambda)}<0$. Since all eigenvalues are positive, they all lie on the right hand side of the vertical asymptote.
For the remaining case, we first have that
$\gamma$ $<1$ (45)
$\gamma^{-1}$ $>1$ (46)
$\gamma^{-1}-\exp(-r(s)/\lambda)$ $>1-\exp(-r(s)/\lambda).$ (47)
Since $\gamma>\exp(r(s)/\lambda)$ and we assume $r(s)<0$, both sides of the inequality are negative. Then,
$\frac{1}{\gamma^{-1}-\exp(-r(s)/\lambda)}$ $<\frac{1}{1-\exp(-r(s)/\lambda)}$ (48)
$\frac{-1}{\gamma^{-1}-\exp(-r(s)/\lambda)}$ $>\frac{1}{\exp(-r(s)/\lambda)-1}.$ (49)
Since all eigenvalues are less than or equal to $\frac{1}{\exp(-r(s)/\lambda)-1}$, they all lie on the left hand side of the vertical asymptote. We have thus shown that when a vertical asymptote exists, it lies outside of the range of eigenvalues, and so the orders of the eigenvectors of the SR and DR are still identical. ∎
@@ B.3 Extension of the DR to State-Action-Dependent Reward
In this subsection, we derive the DR for the case when the reward function depends on both the state and action. The extension of linearly solvable MDPs to the state-action case was first presented by  Ringstrom et al. [37]. However, since (1) the DR was not defined in this work, and (2) there are minor differences in the formulation, derivation, and notation, we present the full derivation below nonetheless. Note, also, that the derivation bears similarity to relative entropy policy search [32], which constrains the KL divergence between an observed data distribution and that generated by the policy.
For this setting, instead of formulating the deviation cost as $\operatorname{KL}\big(p^{\pi}(\cdot|S_{t})\|p^{\pi_{d}}(\cdot|S_{t})\big)$, we formulate the deviation cost from the default policy as $\operatorname{KL}\big(p^{\pi}(\cdot,\cdot|S_{t},A_{t})\|p^{\pi_{d}}(\cdot,\cdot|S_{t},A_{t})\big)$, where $p^{\pi}(\cdot,\cdot|S_{t},A_{t})$ denotes the distribution over the next state-action pairs given the current state-action pair $(S_{t},A_{t})$ and the policy $\pi$. Then, at every time step $t$, the agent receives a reward of $\tilde{r}(S_{t},A_{t})=r(S_{t},A_{t})-\lambda\operatorname{KL}\big(p^{\pi}(\cdot,\cdot|S_{t},A_{t})\|p^{\pi_{d}}(\cdot,\cdot|S_{t},A_{t})\big)$, where $\lambda>0$ determines the relative importance of the deviation cost.